### v2

# Kekulé Superconductivity in Twisted Magic Angle Bilayer Graphene

Abstract
While it has been one of the most important new physics discoveries in the last decade, the nature of superconductivity in the twisted graphene family remains an unsolved problem. Motivated by recent scanning tunneling experiments that report Kekulé ordering in moiré graphene superconductors, we develop a microscopic theory of this superconductivity for the twisted bilayer system. The pairing we find is an intra-valley, finite-momentum pair-density wave (PDW) that intrinsically carries a Kekulé modulation. This state exhibits four salient features: (i) spontaneous breaking of $C_{3}$ rotation symmetry, producing nematic order (ii)with triplet pairing; and (iii) a quasiparticle density of states that evolves from a V-shaped profile to a fully gapped, U-shaped spectrum as the attraction increases which is accompanied by (iv) systematic behavior of the temperature dependent zero bias conductance. These features align with key experimental signatures. We find, as well, that with only modest interaction strengths, the state is near to a BEC-like phase, consistent with the observed extremely short coherence lengths. Taken together, these results identify a microscopic intra-valley Kekulé PDW as a compelling candidate for unconventional superconductivity in the twisted graphene family.
I Introduction
It is hard to overestimate the excitement which has been generated by the discovery [Cao2018a] of “high” temperature superconductivity associated with the flat band regime [Bistritzer2011] in twisted bi- and tri-layer graphene [Nuckolls2024]. Measured so far are transport properties [Polshyn2019, Cao2020], superfluid density[Banerjee2025, Tanaka2025], tunneling[Jiang2019, Oh2021, Xie2019], upper critical fields [Cao2018a, Park2021], thermodynamical [Zondiner2020] as well as other superconducting characteristics[Yankowitz2019, Sharpe2019, Stepanov2019, Saito2019].
Among these, scanning tunneling spectroscopy and microscopy [Nuckolls2023, Kim2023] have led to progress in understanding the nature of the correlated insulators. What is most remarkable is that evidence for an inter-valley coherent (IVC) ordering, called Kekulé was anticipated theoretically [Wagner2022, Kwan2021]. This appears at $\nu=\pm 2$ and $\pm 3$ which coincides with a $\sqrt{3}\times\sqrt{3}$ reconstruction of the atomic scale unit cell. While there has been much theoretical attention [Wagner2022, Kwan2021] paid to the implications for the correlated insulator regimes[Nuckolls2023, Kim2023], a form of Kekulé ordering has also been observed within the superconducting and pseudogap phases in the same interval ($\nu=\pm 2$ to counterpart $\pm 3$.). This, we argue here, is essential to address in order to understand the superconductivity in these twisted graphene, flatband systems.
What has been emphasized in the experimental literature is that (i) the superconductor arises out of a Kekulé order that is qualitatively different from that of the neighboring insulator[Nuckolls2023], (ii) that the superconducting and pseudogap phases look similar to one another with, thus far, no readily distinguishable features in imaging [Nuckolls2023]. (iii) Importantly, the strength of the IVC order, quantified by the normalized Kekulé peak intensity is maximum at filling factors corresponding to those with very strong pairing, as evidenced by the short coherence lengths and the presence of a pseudogap phase[Kim2023].
Figure 1: Cartoon of the twisted bilayer graphene geometry and Cooper Pairs. This serves to contrast intravalley and intervalley pairing. Two graphene Brillouin zones (BZs) are rotated by $\pm\theta$ (gray/red). Two mini-moiré BZs associated with $K/K^{\prime}$ valleys are represented by orange and blue; the $\Gamma/M$-point is shown in the right/left mini-BZ. Intravalley pairing involves pairs within each valley and associated mini-BZ, while intervalley pairing involves two different mini-BZs.
Motivated by this STM data, we now consider intra-valley pairing superconductivity in twisted bilayer graphene (TBG), which is a methodology for implementing superconducting Kekulé order [Roy2010, Tsuchiya2016, yafis2025]. We will see that one can alternatively view this intravalley pairing as a pair density wave (PDW) in which the pairs have non-zero net momentum. In this way, it should be emphasized that “Kekulé superconductivity” as discussed here refers not to a Kekulé pairing mechanism or “glue” but to a Kekulé pairing machinery.
In a qualitative way we can now understand the STM experiments. From (i) it would then appear that one is seeing two distinct Kekulé orders that live in different sectors: in the insulator, a particle-hole Kekulé (bond/charge/valley-coherent) order, and in the superconductor, a particle-particle Kekulé component. From (ii) the observation that pseudogap and superconducting phases show (essentially) the same Kekulé pattern fits naturally with a Kekulé driven pseudogap containing the same preformed pairs as those that become condensed within the superconducting phase. From (iii) that the highest Kekulé peaks are associated with the strongest pairing seems to strongly support the idea that Kekulé ordering and the superconducting pairing machineries are correlated  [^1] To be precise, this correlation should not be viewed as a “Kekulé-driven pairing mechanism”. As shown in arXiv:2506.18996, quite generally, the strongest pairing tends to occur where the non-superconducting ordering (in this case, Kekulé) is weakest, as in a quantum-critical-point scenario and also more generally..
This analysis underlines the importance of systematically addressing, as we do here, the intra-valley pairing case which leads to Kekule superconductivity. This proposal stands in contrast to the majority of current theories, which focus on inter-valley pairing [You2019, Wu2018, Lian2019, Cea2021, Hu2019, Julku2020, Christos2023] (and its variants [Khalaf2021, Gonzalez2019, Po2018]). The simple representation in Figure 1 emphasizes the differences in these two scenarios. One cannot, however, rule out that inter-valley and intra-valley pairing may coexist; they originate from distinct, orthogonal symmetry channels of the underlying microscopic attraction.
We emphasize throughout that this intra-valley pairing scenario leads to important and distinct theoretical consequences from the inter-valley case, some of which are rooted in the structure of the flat-band Bloch wavefunctions. The presence of an effect which we term the “quantum textures” presents a notable contrast. Unlike conventional inter-valley pairing where details about the wavefunction are generally absent from the gap equation, this quantum texture, containing wavefunction information and multi-band effects, enters directly into the intra-valley gap equation. It strongly modulates the behavior of the order parameter. We will see the interplay between this texture and the fundamental symmetries of a single moiré valley dictates the nature of the resulting superconducting order.
II PDW Order and Microscopic Model
A pair density wave superconductor [Agterberg2020] involves pairs with finite center-of-mass momentum, $2\mathbf{Q}$. The order is bond-centered, forming on the AB/BA sublattice bonds, and in the low-filling regime, the pairing momentum $\mathbf{Q}$ is naturally expected to lie near the Dirac point $\bf K$}. The full order parameter is a time-reversal-symmetric combination of condensates from the two valleys (e.g., $\Delta_{2\mathbf{Q}},\Delta_{-2\mathbf{Q}}]$).
[…]
Another crucial constraint on the nature of this PDW superconductivity comes from the fact that STM experiments cannot probe a PDW directly but rather are sensitive to its consequences through the charge sector which must reveal the specific form of Kekulé order. One important consequence of these two-valley PDW states is that they can induce secondary order, involving charge-density waves (CDWs). A CDW arises from the bilinear coupling of the two valley condensates,
$\rho(\mathbf{r})\;\propto\;\Delta_{2\mathbf{Q}}(\mathbf{r})\,\Delta^{*}_{-2\mathbf{Q}}(\mathbf{r})+\text{h.c.},$ (1)
which carries total momentum $4\mathbf{Q}$. On the atomic lattice this reduces to a modulation $\propto\cos(2\mathbf{K}\!\cdot\!\mathbf{r})$ (since $3\mathbf{K}$ is a reciprocal vector), yielding a Kekulé $\sqrt{3}\!\times\!\sqrt{3}$ superlattice, i.e., a tripled unit cell. The slowly varying moiré-scale envelope is set by $\boldsymbol{\delta q}=2\mathbf{Q}-2\mathbf{K}$ and thus varies on the length scale $|\boldsymbol{\delta q}|^{-1}$. Although a full analysis of this secondary order is beyond the scope of this work, this mechanism provides a connection between the proposed PDW and the STM experiments.  [^2] A coexisting intervalley condensate with zero center-of-mass momentum, $\Delta_{0}$, can also couple to the PDW order to induce a secondary CDW: The charge modulation from this second mechanism has a momentum of $2\mathbf{Q}$ and, like the first, behaves as $\cos(2\mathbf{K}\cdot\mathbf{r})$ on the atomic scale. This will then lead to an observed Kekule pattern in the charge channel which can be observed in STM.
Although intra-valley pairing takes place in two different valleys, it is sufficient for our purposes to analyze a single valley separately. That the two valleys contribute independently is well-justified for small twist angles. Our microscopic theory is based on the Bistritzer-MacDonald (BM) continuum model for TBG[Bistritzer2011, PhysRevB.100.035448, PhysRevB.99.035111], supplemented with a short-range attractive interaction. The BM model captures the essential physics of the moiré superlattice through two interlayer tunneling parameters, $w_{AA}$ and $w_{AB}$. A generic translationally invariant pairing interaction can be written as:
$\begin{split}\hat{V}=\sum V_{aa^{\prime}}(\mathbf{q}-\mathbf{q}^{\prime})\,&\psi^{\dagger}_{a,\sigma}(\mathbf{k}_{+})\psi^{\dagger}_{a^{\prime},\sigma^{\prime}}(\mathbf{k}_{-})\,\\ &\times\psi_{a^{\prime},\sigma^{\prime}}(\mathbf{k}^{\prime}_{-})\psi_{a,\sigma}(\mathbf{k}^{\prime}_{+}),\end{split}$ (2)
where the momenta are defined as $\mathbf{k}_{\pm}=\mathbf{k}\pm\mathbf{q}$ and $\mathbf{k}^{\prime}_{\pm}=\mathbf{k}^{\prime}\pm\mathbf{q}^{\prime}$. The summation is over all momenta ($\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\mathbf{q}^{\prime}$), spin indices ($\sigma,\sigma^{\prime}$), and layer/sublattice indices ($a,a^{\prime}$).
We require pairing interactions to respect all symmetries of the single valley BM model, and involve two different sub-lattices, as we consider AB-bonding order. Among the possible interactions that satisfy these criteria, a natural choice [Roy2010, Tsuchiya2016] is a nearest-neighbor (NN) attraction within each layer. This is physically motivated by the fact that the interlayer distance is generally larger than the intralayer nearest-neighbor distance.
For definiteness, the numerical results and figures presented throughout this work are based on this NN interaction model. As we will show, the essential properties of the resulting superconducting order originate from the symmetry properties of the BM model. Importantly they are viewed as generally generic and, thus, independent of the details of the assumed short-range attractive interactions.
We begin by utilizing the spatial $C_{3}$ rotational symmetry of the lattice to classify the possible superconducting pairing channels. The $C_{3}$ point group has three one-dimensional irreducible representations, which we label as $A$, $E_{1}$, and $E_{2}$. An eigenstate belonging to one of these irreducible representations acquires a phase of $\zeta^{l},\,l=0,1,2$ under a $C_{3}$ rotation, where $\zeta=\exp(i2\pi/3)$. The interaction potential can be decomposed into these eigenstates. For the nearest neighbor model, we have the decomposition for each layer:
$V(\mathbf{q}-\mathbf{q}^{\prime})=\sum_{l=0}^{2}g_{l}f_{l}(\mathbf{q})f_{l}^{*}(\mathbf{q}^{\prime}).$ (3)
Here $f_{l}(\mathbf{q})=3^{-1}\sum_{j=0}^{2}\zeta^{jl}\,\exp{-i\mathbf{q}\cdot\boldsymbol{\delta}_{j}}$ , and $\boldsymbol{\delta}_{j}$ represents the nearest-neighbor vectors in the honeycomb lattice.
As shown in Figure 1, for intra-valley pairing, Cooper pairs are formed by electrons within the same moiré valley As the mini-Brillouin zone is little, this implies that the interaction is dominated by small momentum transfers, $\mathbf{q}$. Consequently, the phase factor in each $f$ is small [^3] Rigorously, the momentum transfer involved is $\mathbf{q}+\mathbf{G}_{j}$, where $\mathbf{q}$ lies within the first mini-Brillouin Zone and $\mathbf{G}_{j}$ is a moiré reciprocal lattice vector. In the Bistritzer-MacDonald model, $|\mathbf{G}_{j}|$ can be up to three or four times the magnitude of the primitive moiré reciprocal lattice vectors. Nevertheless, the total momentum $|\mathbf{q}+\mathbf{G}_{j}|$ remains significantly smaller than any momentum associated with the atomic scale. $|\mathbf{q}\cdot\boldsymbol{\delta}_{j}|\ll 1$. In this limit, the form factor for the $A$-channel ($l=0$) approaches $f_{0}(\mathbf{q})\to 1$, while the form factors for the chiral channels ($l=1,2$) vanish: $f_{1,2}(\mathbf{q})\to 0$. Thus the dominant superconducting channel corresponds to $A$  [^4] This conclusion is general and applies to a large class of short-range interaction, as the argument holds if the set of nearest-neighbor vectors, $\boldsymbol{\delta}_{j}$, is replaced by any short-range atomic vector and its $C_{3}$ rotational partners. .
Figure 3: The figure shows the grand-canonical thermodynamic potential $\Omega$ of the singlet and triplet pairing states as a function of interaction strength $V\equiv V_{0}(q=0)$. Results are computed at near-zero temperature $T=0.001~\mathrm{meV}$ and chemical potential $\mu=-0.59~\mathrm{meV}$. The plotted quantity is $\Omega-\Omega_{N}$, measured relative to the normal-state value $\Omega_{N}$. The inset shows the corresponding superconducting order-parameter magnitude $|\Delta|$ for the two states. For the interaction strengths shown, the triplet state has a lower $\Omega$ than the singlet state, indicating that triplet pairing is thermodynamically favored.
III Order Parameter and Quantum Textures
For this case which we consider throughout the paper, the pairing gap parameter satisfies:
${\Delta}_{aa^{\prime},\sigma\sigma^{\prime}}(\mathbf{Q})=-\sum_{\mathbf{q}}V_{0}(\mathbf{q})\,\langle\hat{\psi}_{a^{\prime},\sigma^{\prime}}(\mathbf{Q}_{-})\hat{\psi}_{a,\sigma}(\mathbf{Q}_{+})\rangle.$ (4)
Here $V_{0}(\mathbf{q})=g_{0}f^{*}_{0}(\mathbf{q})$ and $\mathbf{Q}_{\pm}=\mathbf{Q}\pm\mathbf{q}$. In the NN model, we have the layer/sublattice indices $a=LA$ and $a^{\prime}=LB$. This is consistent with the physics of a condensate which forms on the AB sublattice bonds within each layer. Furthermore, the condensates in the two layers are related by mirror symmetry such that they are equal in magnitude.
To analyze the pairing further, we rewrite the order parameter in the spectral (or band) basis. The transformation from the plane-wave basis (operators $\hat{\psi}$) to the spectral basis (operators $c$) requires the introduction of the Bloch wavefunctions, $u_{\mathbf{G},a;n}(\mathbf{k})$:
$\hat{\psi}_{a,\sigma}(\mathbf{k}+\mathbf{G})=\sum_{n}u_{\mathbf{G},a;n}(\mathbf{k})\,\hat{c}_{n,\sigma}(\mathbf{k})$ (5)
This leads to the important definition of a (particle-particle) form factor, $\Lambda_{mn,aa^{\prime}}(\mathbf{k},\mathbf{q})$, which projects the interaction onto the particle-particle channel in this bandstructure basis:
$\Lambda_{mn,aa^{\prime}}=\sum_{\mathbf{G}}f^{*}_{0}(\mathbf{q}+\mathbf{G})\,u_{\mathbf{G},a;m}(\mathbf{Q}_{+})\,u_{-\mathbf{G},a^{\prime};n}(\mathbf{Q}_{-}).$ (6)
Using this form factor, Eq. 4 can be rewritten in terms of:
${\Delta}_{aa^{\prime},\sigma\sigma^{\prime}}=g_{0}\sum\Lambda_{mn}(\mathbf{Q},\mathbf{q})\,\langle\,\hat{c}_{m,\sigma}(\mathbf{Q}_{+})\hat{c}_{n,\sigma^{\prime}}(\mathbf{Q}_{-})\rangle.$ (7)
Here the summation is over the band index $m/n$ and momentum $\bf q$} in the mini-BZ. We emphasize that in the form factor of Eq. 6, only direct products of Bloch wavefunctions appear. This is to be contrasted with their inner products which would appear in the particle-hole sector. Time-reversal symmetry relates opposite valleys and is absent within a single moiré valley. Consequently, the diagonal form factor $\Lambda_{nn}(\mathbf{Q},\mathbf{q}=0)$ generally deviates from unity, and the off-diagonal components $\Lambda_{mn}$ acquire sizeable magnitudes that are central to the pairing machinery. This structure, which reflects the unique quantum texture of the intra-valley sector, is addressed more explicitly in Fig. 2. This is a key distinction between intra-valley pairing and BCS-like inter-valley pairing (or the correlated Kekulé insulator), where the analogue diagonal form factors are close to unity.
In our numerical simulations, we perform a truncation of those remote bands in the vicinity of the flat-bands. This analysis confirms that a two-isolated- flat- band approximation is adequate, largely because of the sizeable separation $E_{\text{band}}$ between the flat and remote bands. Here we have that the bandwidth $W$ (which is very small in the absence of strain) satisfies $\Delta/E_{\text{band}}\ll W/\Delta$. More precision can be obtained when more bands are added, as is discussed in the Methods section.
The pairing physics is governed by a form factor matrix, $\Lambda_{mn}$, where the indices $m,n$ span the band basis. We refer to this matrix as a “quantum texture” which plays a central role in our analysis as it reflects the momentum and band space structure of the pairing interaction, thus determining the fundamental properties of the condensate.
III.1 Establishing the Thermodynamically Stable Phase
For a given pairing momentum $\bf Q$}, one can solve the gap equation self-consistently to find possible solutions. Because not all solutions represent thermodynamically stable states it is necessary to investigate their stability by establishing that they correspond to a global minimum of the grand-canonical potential $\Omega$ at a given chemical potential $\mu$.
In the PDW state under consideration, the thermodynamic potential is a function $\Omega(\mu,{\bf Q},\Delta_{L},{\bf S})$, where we label the layer index $L=1,2$ and $\mathbf{S}$ represents the pair spin. There are two branches of solutions satisfying the gap equation: $\Delta_{1}=\Delta_{2}$ and $\Delta_{1}=-\Delta_{2}$. Both branches respect mirror symmetry. The first branch, which is the physical solution, corresponds to a global minimum for a fixed $\bf Q$} and $\mu$, while the second branch represents a saddle point. Thus, throughout the paper we use $\Delta\equiv\Delta_{1}$ to denote the order-parameter magnitude.
In this stable ground state with $\Delta_{1}=\Delta_{2}$, $\bf Q$} is a variational parameter which is constrained to be within the mini-BZ  [^5] The zero-current condition is automatically satisfied once the time-reversal symmetric partner at the other valley is included. This differs from the known FF state, where $Q$ must be solved self-consistently to satisfy the zero-current condition. Important here also is that a conventional FF state is generally associated with a saddle point (see Wang et al, Phys. Rev. B 97, 134513), which is circumvented in multilayer systems where there are more degrees of freedom to enable stability.. For a given $\mu$ and spin configuration (singlet or triplet), the potential $\Omega$ is minimal at ${\bf Q}=M$ in the mini-BZ. This indicates that the slow moiré-scale modulation observed in STM data must be determined by the momentum $M$ relative to Dirac points.
We note that the relative magnitude of $\Omega$ for the singlet and triplet states depends on the pairing momentum $\bf Q$}. Importantly, when $\bf Q$} is chosen to be the $M$-point, the triplet state has a lower $\Omega$ than the singlet, as is consistent with what is observed experimentally [Cao2021]. This is shown in Fig. 3, where it should be clear that the energy scales associated with the strength of the needed attraction are of the order of meV. This should be quite accessible and it is notably smaller than estimates [Julku2020] of the order of $0.1$ eV, based on inter-valley pairing. This can be seen in more detail in Fig. 4.
The difference between the singlet and triplet pairings [^6] Since the Hamiltonian preserves full spin rotation symmetry, we select the triplet channel with $S_{z}=0$ without loss of generality. is encoded in the structure of the form factors:
$\Lambda^{s/t}(\mathbf{Q},\mathbf{q})=\frac{1}{2}\left[\Lambda(\mathbf{Q},\mathbf{q})\pm\Lambda^{\mathsf{T}}(\mathbf{Q},-\mathbf{q})\right].$ (8)
Here, $s$ and $t$ denote singlet and triplet, respectively, and the transpose acts on the band indices. Due to $C_{2}\mathcal{T}$ symmetry, the diagonal component of $\Lambda^{t}$ (triplet channel) is strongly suppressed compared to the singlet channel [^7] The $C_{2}\mathcal{T}$ symmetry enforces the condition $\Lambda^{t}_{nn}(\mathbf{Q},\mathbf{q}=\mathbf{0})=0$.. However, there are two active flat bands near the chemical potential. As in Ref. [Christos2023] which considered a different pairing configuration, the off-diagonal pairing plays a crucial additional role here regarding the competition between singlet and triplet states which depend on $\bf Q$}. To summarize, the stable order is characterized by two equal layer condensates, pairing momentum at $M$, and spin-triplet pairing.
Figure 4: Order-parameter amplitude $|\Delta|$ (left axis) and chemical potential $\mu$ (right axis) versus coupling strength $V$. The dashed line marks the flat-band bottom. For attractive $V\approx-1.7\,\mathrm{meV}$ ($\Delta\simeq 0.73$meV), the system crosses into the BEC-like regime.
III.2 Nematic Nature of Superconducting order
The superconducting order order under consideration is generically nematic, spontaneously breaking the three-fold rotational symmetry $C_{3}$. This can be directly observed in the quantum texture in Fig. 2(c),(d). This anisotropy does not originate through a pre-determined pairing interaction, but rather the $C_{3}$ rotational symmetry is broken as a consequence of condensation.
Note that this is different from the inter-valley case, where nematicity will emerge after selecting a particular combination of the $E$-doublet channels $f_{1}/f_{2}$. A rigorous proof for the intra-valley case is provided in the Methods section. This proof establishes the fact that in the BM bandstructure, there is a $C_{3}$ rotational symmetry about the $\Gamma$ point of the mini-BZ, but the condensed state is not invariant under a $C_{3}$ rotation. Hence there is spontaneous symmetry breaking (SSB) of $C_{3}$ symmetry.
In the Methods section we show how this happens by noting that a $C_{3}$ rotation essentially causes a momentum boost to the Bloch wavefunction. Equivalently, one can see that there is a unitary operator satisfying
$U_{C_{3}}(\mathbf{r})H(\mathbf{r})U_{C_{3}}(\mathbf{r})^{\dagger}=H(C_{3}^{-1}\mathbf{r}),$ (9)
where $U_{C_{3}}(\mathbf{r})\sim\exp(i\mathbf{p}_{L}\cdot\mathbf{r})$ and $\mathbf{p}_{L}$ is a layer-dependent momentum. This boost necessarily changes the pairing operator $\hat{\Delta}(\mathbf{r})$ and shifts the momentum of the condensate, $\mathbf{Q}$, to $\mathbf{Q}+\mathbf{p}_{L}$.
Figure 5: (a) Density of states (DOS) versus energy $\omega$ (meV) for three different gap amplitudes, with curves labeled by $|\Delta|\simeq 0.49,0.36,$ and $0.32$ meV. As the order parameter is reduced, the bottom of the characteristic U-shaped gap narrows towards a V-shape. Notably, a finite zero-bias conductance emerges in the latter case (b) The lowest positive Bogoliubov-de Gennes eigenvalue $E_{1}(M,\mathbf{k})$ plotted across the mini-Brillouin zone for $|\Delta|=0.49$ meV. At this amplitude (still within the BCS-like regime), the spectrum remains fully gapped, corresponding to the U-shaped DOS. (c). At a smaller amplitude of $|\Delta|=0.32$ meV, a small Bogoliubov Fermi surface appears and the (V-shaped) DOS spectrum becomes gapless.
IV U to V-like transitions in tunneling
In the trilayer case, where Kekulé superconductivity is similarly observed [Kim2023] the evolution of the conductance with doping is of particular interest. The tunneling spectrum which reflects the quasi-particle density of states [Kim2022], has a V-like shape (often with a “zero bias conductance”) when filling is in the regime closer to $\nu=-3$, while it transitions to a U-shaped spectrum, when $\nu$ is nearer 0.2-0.2. It is generally believed that these spectra, then, reflect a transition from a gapped superconductor to a nodal superconductor. Important to this analysis, one can associate the U regime with a shorter Landau Ginsberg coherence length which reflects [Chen2024] stronger attractive interactions. In this context, while originally this U transition was speculated to relate to a transition all the way to a BEC regime, this may have been refuted [Van2025].
Although these experiments address the tri-layer case, it is useful here to study the density of states (DOS), using the parameter set corresponding to Fig. 2. As shown in Fig. 5(a), for a superconducting order parameter of $|\Delta|\geq 0.4$ meV, the DOS exhibits a clear U-shape, indicating that the quasiparticle excitations are fully gapped.
As the order parameter is reduced, the flat bottom of this U-shaped gap narrows, eventually transitioning into a smoothed V-shape with a finite DOS at zero frequency. We find this corresponds to the fact that there are touching points leading to a very small Bogoliubov Fermi surface contained within the quasi-particle dispersions for the “V case” and missing for the “U case”, as can be seen in Fig. 5(b) and Fig. 5(c)  [^8] Although their Bogoliubov Fermi surfaces are larger, there is some similarity here to work from Christos et al, Nat. Commun. 14, 7134 and Agterberg et al, Phys. Rev. Lett. 118, 127001. We stress here that there is generally a finite conductance at zero energy (zero bias conductance) when the V shape is present. This constitutes a prediction from the present theory.
To address this in more detail, it is useful to consider the Bogoliubov dispersion in the two-flat-band limit:
$E(M,{\bf q})=\delta\epsilon({M,\bf q})\pm\sqrt{\bar{\epsilon}({M,\bf q})^{2}+\Delta^{2}\varphi({M,\bf q})^{2}}$ (10)
where $\delta\epsilon({M,\bf q})=[\epsilon_{1}(M+{\bf q})-\epsilon_{2}(M-{\bf q})]/2$ and $\bar{\epsilon}({M,\bf q})=[\epsilon_{1}(M+{\bf q})+\epsilon_{2}(M-{\bf q})]/2$ are derived from the dispersion of two flat bands, $\epsilon_{1,2}({M,\bf q})$. Here $\varphi(M,q)=\sum_{L}\Lambda^{t}_{12,L}(M,q)$ represents the strength of off-diagonal pairing [^9] From Fig. 2, we observe that the off-diagonal components of $\Lambda$ are significantly larger than the diagonal ones. Consequently, we retain only the off-diagonal pairing..
From this expression, two regimes emerge. When $\Delta$ is large enough such that the condition $|\Delta\varphi(M,{\bf q})|>|\delta\epsilon(M,{\bf q})|$ holds for all $\bf q$}, the dispersion is necessarily gapped, leading to the U-shaped DOS. Conversely, when $\Delta$ is small enough that $|\Delta\varphi(M,{\bf q})|\leq|\delta\epsilon(M,{\bf q})|$ for some $\bf q$}, gapless points can develop. This results in a V-shaped DOS generally associated with a small Bogoliubov Fermi surface, reflecting the sharpness of the V. We anticipate the existence of a critical value, $\Delta_{c}$, that marks a transition from a fully gapped to a gapless superconducting state.
V Zero-bias Conductance
Of interest is the behavior of the zero-bias conductance in these tunneling spectra. The fact that two different groups [Park2025, Kim2022] have arrived at quite similar plots suggests that this may be an intrinsic phenomenon. What is important here is that the zero bias conductance, in principle, contains information about the excitation spectrum of the superconductor. Moreover, there should be some distinction between the U and V shaped cases which will be important to assess in future, although the experiments thus far have been in the V-shaped regime.
We present a plot of the zero bias conductance in this regime, as a function of temperature which is compared with the temperature dependence of order parameter $\Delta$. This is shown in Fig. 6 and is inspired by a related plot (see their Fig 3e) for the moiré materials in Ref. [Park2025]. In the Methods section we discuss the U-shaped counterpart.
One can speculate that here the V-shaped tunneling spectra may be qualitatively different from the classic $d$-wave behavior in the cuprates [Pushp2009] where a plot $dI/dV\propto V$ vanishes exactly at zero bias. Throughout the cuprate family, however, there are some indications of a finite zero bias conductance but these are generally understood [Reber2013] as associated with a Dynes lifetime parameter in the tunneling density of states. We argue that what is different about the moiré systems is that the finite zero-bias conductance appears order-parameter controlled, not merely lifetime-broadened. This suggests a superconducting state whose excitation spectrum remains partially gapless due to intrinsic features of the paired state itself, as illustrated in Fig. 6. Here the theory for the graphene case reflects the presence of a Bogoliubov Fermi surface. Given the consistency across a few fillings  [^10] Private Communications with Hyunjin Kim and across groups [Park2025, Kim2022], we attribute the finite zero-bias conductance to an intrinsic Bogoliubov Fermi surface rather than to extrinsic, random Dynes-like broadening.
Figure 6: Density of states and order parameter versus temperature, computed at the interaction strength V=0.8V=-0.8meV. As temperature increases, the order parameter decreases while the DOS increases. This behavior is consistent with the presence of a Bogoliubov Fermi surface and can be seen to be related to Figure 3e in Ref. [Park2025] (see text), which addresses the zero-bias conductance.
VI Conclusions
In addition to these zero bias predictions, we have three central results concerning the twisted graphene family in this paper which we argue have some support from experiment: these are nematic order [caoNematic], U to V transitions [Kim2022] and triplet pairing [Cao2021]. All these results are generic consequences of Kekulé superconducting order and symmetry properties of the single moiré valley (BM) model. At the heart of our paper is the assumption of intra-valley (Kekulé) pairing in which the major support comes from an observed correlation [Kim2023] in these moiré systems between the highest Kekulé peaks which are associated with the strongest pairing regime.
We argue against the interpretation that Kekulé order acts as a pairing ”glue” since we know in superconductors which have a non-phononic, bosonic (e.g., magnetic) pairing mechanism, the superconductivity is strongest when the non-superconducting, alternative ordering is weakest[zq2025]. Rather Kekulé order seems to be intertwined with the pairing machinery as distinct from the pairing mechanism. This order then leads to a PDW-like description of the superconductivity[Tsuchiya2016, Roy2010, PhysRevLett.114.237001, wang2025].
We have, throughout this paper, made few distinctions between the bi- and tri-layer moiré graphene superconductors. We believe, in line with most of the experimental phenomenology, that many of the general features discussed here apply to both; we focus almost exclusively on the physics of the two flat bands and the related symmetries. There is an interesting distinction however which is that the attractive interactions (as measured by the large size of the pairing gap and the very short Landau Ginsberg coherence lengths) may not be strong enough in the bilayer case to reveal a U shaped tunneling spectra. The V shape features should remain and it would be important to see if there is consistency with the theoretically expected trends which associate this regime with a finite zero bias conductance.
In summary, it is our contention in this paper that this Kekulé/PDW form of superconductivity is an essential ingredient which must be included in order to ultimately understand superconductivity in the twisted graphene family.
Appendix A Methods
A.1 Bistritzer–MacDonald Model and Symmetries
We briefly review the Bistritzer–MacDonald (BM) model of twisted bilayer graphene (TBG) and its symmetries. When two layers are stacked with a relative twist angle $\theta$, a fundamental momentum scale emerges,
$k_{\theta}=2K\sin\frac{\theta}{2},\qquad K=\frac{4\pi}{3\sqrt{3}a},$ (11)
where $a$ is the graphene lattice constant. A convenient choice of momentum difference between the two Dirac points is
$\mathbf{q}_{1}=\mathbf{K}_{\text{top}}-\mathbf{K}_{\text{bot}}=(0,1)\,k_{\theta}.$ (12)
The remaining moiré reciprocal vectors can be defined as $\mathbf{b}_{1,2}=\mathbf{q}_{2,3}-\mathbf{q}_{1}$, where $\mathbf{q}_{2,3}$ are obtained by counterclockwise rotations of $\mathbf{q}_{1}$ by $120^{\circ}$.
The BM Hamiltonian describes the continuum theory near the two Dirac cones,
$H=v\Big(-i\nabla-L_{z}\mathbf{q}_{1}/2+\mathbf{M}\Big)\cdot\boldsymbol{\sigma}_{\theta/2}+\big[T(\mathbf{r})L^{+}+\text{h.c.}\big],$ (13)
where $v$ is the Fermi velocity of monolayer graphene, $\mathbf{M}$ denotes the $M$-point of the mini Brillouin zone, $L$ and $\sigma$ act on layer and sublattice indices, and $\boldsymbol{\sigma}_{\pm\theta/2}$ denote Pauli matrices rotated by angle $\pm\theta/2$. The tunneling matrix $T(\mathbf{r})$ is parameterized by two amplitudes $w_{AA}$ and $w_{AB}$,
$T(\mathbf{r})=w_{AA}\,\alpha_{0}(\mathbf{r})+w_{AB}\,\sigma^{+}\alpha_{1}(\mathbf{r})+w_{AB}\,\sigma^{-}\alpha_{2}(\mathbf{r}),$ (14)
with spatial form factors
$\alpha_{m}(\mathbf{r})=\sum_{j=0}^{2}w^{mj}\,e^{-i\mathbf{b}_{j}\cdot\mathbf{r}},\qquad b_{0}={\bf 0},\quad w=e^{i2\pi/3}.$ (15)
The BM Hamiltonian respects four fundamental symmetries: moiré translations, threefold rotations $C_{3}$, mirror reflection $M_{y}$, and the combined twofold rotation with time reversal $C_{2}\mathcal{T}$. For a spatial transformation $R$, the Hamiltonian transforms as
$H^{\prime}(\mathbf{x})=H(R^{-1}\mathbf{x}),$ (16)
and invariance under the symmetry requires
$H^{\prime}(\mathbf{x})=U(\mathbf{x})H(\mathbf{x})U^{\dagger}(\mathbf{x}),$ (17)
with $U(\mathbf{x})$ a unitary operator. Explicitly, the mirror symmetry and threefold rotation[PhysRevB.99.035111] are represented by $U_{M_{y}}=\sigma_{x}L_{x}$ and
$U_{C_{3}}(\mathbf{x})=\exp\!\left(\tfrac{2\pi i}{3}\,\hat{z}\cdot\boldsymbol{\sigma}\right)\exp\!\left[i\Big(2\mathbf{M}+(L_{z}+1)\mathbf{b}_{1}/2\Big)\cdot\mathbf{x}\right].$ (18)
For the antiunitary $C_{2}\mathcal{T}$ operation, one has
$H^{\prime}(\mathbf{x})=H^{*}(-\mathbf{x}),\qquad H^{\prime}(\mathbf{x})=\sigma_{x}H(\mathbf{x})\sigma_{x},$ (19)
which guarantees invariance under the combined symmetry.
A.2 Flat band Approximation
We find that $\Lambda$ is highly nonlocal in the band index: sizeable matrix elements $|\Lambda_{mn}(\mathbf{q})|$ persist even for $|m-n|\gg 1$. Thus it is necessary to investigate the coupling between the flat bands and remote bands. Experimental data and numerical simulations consistently indicate a significant separation of energy scales. The superconducting gap is typically on the order of $\Delta\sim 1~\mathrm{meV}$, whereas the single-particle gap separating the flat and remote bands in the BM model is much larger, at $E_{\text{band}}\sim 50~\mathrm{meV}$. A Schrieffer–Wolff transformation can be employed to integrate out the remote bands:
$\tilde{H}_{LL}\simeq H_{LL}+H_{LH}\,(-H_{HH})^{-1}H_{HL}.$ (20)
Here, $H_{LL}$ is the Hamiltonian in the flat-band subspace, $H_{LH}$ couples the flat and remote bands, and $H_{HH}$ is the Hamiltonian within the remote-band subspace. After the transformation, the new Hamiltonian $\tilde{H}_{LL}$ acquires a correction of order $\Delta^{2}/E_{\text{band}}$. When this energy is much smaller than the bandwidth $W$ of a single flat band, the isolated-flat-band description of intra-valley superconductivity is well-justified, provided the bands are not perfectly flat ($W\!\to\!0$) or $\Delta$ sufficiently large so that remote-band effects must be included.
In numerical simulations that incorporate remote bands into the gap equation, we find that including only the two flat bands yields reasonable results, although adding more bands enhances precision. In practice, we retain $10$ bands.
Figure 7: U-shaped density of states (DOS) from two-flat-band superconductivity. The figure is plotted at $|\Delta|\simeq 0.63~\mathrm{meV}$ ($V\simeq-1.5$ meV). The dip feature around $\omega\sim$ 1meV originates from the band touching of two BM flat bands at $K/K^{\prime}$ points.
Figure 8: Quasi-particle density of states (reflecting zero bias conductance) and order parameter versus temperature, computed at the interaction strength $V\simeq-1.17$meV. This is in the U-shaped tunneling regime and should be contrasted with Fig. 6 in the main text which is for the V-shaped case with weaker attractive interaction. As temperature increases, the density of states remains gapped until the order parameter $\Delta$ drops sufficiently and then the system crosses over to a more V-like spectrum associated with a Bogoliubov Fermi surface.
Appendix B U/V-transitions in the DOS
As discussed in the main text, the U/V transition can be understood through the analysis presented in Eq. (10). For a typical order parameter, say $\Delta\sim 0.3$ meV, the associated Bogoliubov Fermi surface is small, as depicted in Fig. 5. This implies that the “V” shape has a rather sharp minimum.
Here, we present the full plot of the density of states (DOS) within the U-shaped regime in Fig. 7. While the low-energy behavior has already been presented, substructure corresponding to dips are observable around $\omega\sim 1$ meV. These dips arise from the structure of two flat bands that touch at the $K$ and $K^{\prime}$ points. The particle-hole asymmetry in the DOS is a result of our choice of filling which is taken to be $3/8$ of the lower flat band (equivalent to $\nu=-2.5$ filling in experiments).
In Fig. 8 we plot the effective zero bias conductance as well as the order parameter for a strong attraction case where the tunneling curve assumes a U-shape. This represents a prediction as, thus far, there are no published curves for this zero bias conductance. In contrast to Fig. 6 at weaker attraction, here, as temperature increases the zero bias density of states remains gapped until crossing over to the behavior associated with the V-shaped regime where a Bogoliubov Fermi surface is present.
B.1 Proof of $C_{3}$-SSB
Here we provide a proof that the intravalley superconducting order breaks $C_{3}$-symmetry. In the BM model, there is a $C_{3}$ symmetry about $\Gamma$. A counterclockwise $C_{3}$ rotation is represented by the unitary operator in Eq. 18.
The action of $U_{C_{3}}$ on the pairing operator $\hat{\Delta}(\mathbf{r})$ induces a layer-dependent momentum boost:
$\hat{\Delta}_{L,\sigma\sigma^{\prime}}(\mathbf{r})\;\longmapsto\;e^{\,2i\,\mathbf{p}_{L}\!\cdot\mathbf{r}}\;\hat{\Delta}_{L,\sigma\sigma^{\prime}}(\mathbf{r}).$ (21)
Here $\hat{\Delta}(\mathbf{r})$ denotes the operator corresponding to Eq. (4), and only the layer index $L$ is preserved for simplicity, and $\mathbf{p}_{L}=2\mathbf{M}$ for the bottom layer while $\mathbf{p}_{L}=2\mathbf{M}+\mathbf{b}_{1}$ for the top layer. This phase cannot be removed by a global $\mathrm{U}(1)$ gauge choice: it is a layer-dependent linear function of $\mathbf{r}$ (a boost), so gauging it away in one layer necessarily induces an incompatible shift in the other.
First consider condensation at the highest-symmetry $\Gamma$ point of the model and take this as the reference momentum, i.e., the pairing momentum $\mathbf{Q}=0$. Here we examine the symmetry of mean-field Hamiltonian, i.e., fixing condensate and transforming fermions [^11] An equivalent approach is to show that the condensate is not invariant under $C_{3}$. The proof is similar.. Since the BM-model term is invariant under a $C_{3}$ rotation, we only need to examine the mean-field term:
$\sum_{\mathbf{q},m,n}\Delta_{L,\sigma\sigma^{\prime}}\,\Lambda^{*}_{mn,L}(\mathbf{0},\mathbf{q})\,\hat{c}^{\dagger}_{\sigma,m}(\mathbf{q})\,\hat{c}^{\dagger}_{\sigma^{\prime},n}(-\mathbf{q}).$ (22)
A $C_{3}$ rotation acts as
$\Lambda^{*}_{mn,L}(\mathbf{0},\mathbf{q})\;\longmapsto\;\Lambda^{*}_{mn,L}(-\mathbf{p}_{L},\mathbf{q}),$ (23)
which shifts the pairs’ center-of-mass momentum by $\mathbf{p}_{L}$. Recalling the definition in Eq. 6, the Bloch-amplitude product $u_{\mathbf{G}}u_{-\mathbf{G}}$ in the form factor becomes $u_{\mathbf{G}-\mathbf{p}_{L}}u_{-\mathbf{G}-\mathbf{p}_{L}}$, thereby modifying the form factor. Hence the Bogoliubov spectrum is not invariant under $C_{3}$, which is consistent with the form factor presented in Fig. 2.
If the condensate forms at a finite momentum $\mathbf{k}\neq\Gamma$, $C_{3}$ is likewise broken: $\mathbf{k}$ itself rotates under $C_{3}$, and the shift in Eq. (23) still applies. This completes the proof that superconducting order generically breaks threefold rotational symmetry. Importantly, the argument does not rely on any particular form of the attractive interaction; it uses only that the superconducting order is intra-valley.
Appendix C Data availability
The data analyzed in the current study are available from the author Ke Wang on reasonable request.
Appendix D Code availability
The codes used for the current study are available from the author Ke Wang on reasonable request.
Appendix E Acknowledgement
We thank Kevin Nuckolls, Hyunjin Kim, Ivar Martin, Bitan Roy, Stevan Nadj-Perge, Senthil Todadri, Zhiqiang Wang, Ananth Malladi for helpful discussions. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.
Appendix F Author Contributions
K.L. conceived and supervised the project. K.W. performed the computations. K.W. contributed to the acquisition of the data and preparation of figures. All authors have contributed to the interpretation of the data and the drafting as well as the revision of the manuscript.
Appendix G Competing Interests
The authors declare no competing interests.
Appendix H ADDITIONAL INFORMATION
Correspondence and requests for materials should be addressed to the authors K. Wang and K. Levin.